OpenStudy (anonymous):

Solve the equation: x2 - 14x + 85 = 0

OpenStudy (anonymous):

1) Find two numbers that add up to -14 and multiply up to 85 but it won't work in this case so you will have to use the quadratic function \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] 2) a=1, b=-14, and c=85 \[\frac{ 14\pm \sqrt{196-340} }{ 2 }\] \[\frac{ 14\pm \sqrt{-144} }{ 2}\] (you have a negative in the radical and you can use (i)^2 which represents -1) \[\frac{ 14\pm 12\iota }{ 2 }\] (you can simplifly by dividing 14 by 2 and 2 by 2) \[7\pm12\iota \] (this is your final answer) Hope this helps :D